OpenStudy (anonymous):

Given f(x) = the quantity of 4x plus 1, divided by 3 solve for f-1(3).

OpenStudy (johnweldon1993):

\[\large f(x) = \frac{4x + 1}{3}\] We want the inverse of this...so first we make f(x) = y ...and then we switch the positions of the 'x' and 'y' \[\large x = \frac{4y + 1}{3}\] Now we just need to solve for 'y' again... Any idea how to start that?

OpenStudy (anonymous):

no clue math is not my subject at all I struggle

OpenStudy (johnweldon1993):

No problem :) Alright so we have \[\large x = \frac{4y + 1}{3}\] We want to use algebra to solve for 'y' agan...so lets begin by multiplying both sides of this equation by 3...*because it will cancel out the '3' on the right side* \[\large 3x = 4y + 1\] See what I did there?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

Alright...now we want to isolate the 4y...so lets subtract 1from both sides of the equation *to cancel it again on the right side* \[\large 3x - 1 = 4y\] The last step...is to divide everything by 4...this will cancel out the 4y on the ight leaving us with just 'y' \[\large y = \frac{3x - 1}{4}\] This is our inverse...now we just need to evaluate it when x = 3

OpenStudy (anonymous):

Okay that makes sense

OpenStudy (johnweldon1993):

Alright..so then we have \[\large y =\frac{3x - 1}{4}\] \[\large y = \frac{3(3) - 1}{4}\] \[\large y = \frac{9 - 1}{4}\] \[\large \frac{8}{4}\] \[\large y = \space?\]

OpenStudy (anonymous):

y==2

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (anonymous):

Thank you that makes much more sense now!

OpenStudy (johnweldon1993):

Anytime :)